Hi! I was wondering if someone would be able to help me with this problem? I have been having a hard time grasping the graphs in this chapter.
Thank you!!!! :)
Write the slope-intercept equation for the line containing the given pair of points.


(6,8) and (3,5)

Answer by rfer(16322)

m=(5-8)/(3-6)
m=-3/-3
m=1
=============
y=mx+b
5=1*3+b
b=2
=============
y=x+2